40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 2 18:35:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
    Reply  Tue Feb 2 18:54:02 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
       Reply  Wed Feb 3 23:23:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table KnifeEdgeScans_x.pdfBeamscan_x.pdf
          Reply  Thu Feb 4 09:16:46 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
             Reply  Fri Feb 5 00:23:01 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table Beamscan_setup.pdfBeamscan_x.pdfBeamscan_y.pdfZscan.pdf
                Reply  Fri Feb 5 16:50:24 2016, gautam, Update, Green Locking, First pass at mode-matching Modematch_AUXx.pdfNewSetUp.pngModematch_AUXx_2.pdfXendModeMatch.m.zip
                   Reply  Mon Feb 8 15:36:37 2016, gautam, Update, Green Locking, Alternative mode-matching scheme Modematch_alternative.pdf
                      Reply  Tue Feb 9 04:37:10 2016, ericq, Update, Green Locking, Laser swap initiated 
                         Reply  Tue Feb 9 11:49:47 2016, gautam, Update, Green Locking, Laser swap initiated Innolight_beamProfile.pdf
                            Reply  Tue Feb 9 19:15:36 2016, gautam, Update, Green Locking, Laser swap - updates Modematch_X.pdfXendModeMatch.m.zip
                               Reply  Wed Feb 10 17:57:15 2016, gautam, Update, Green Locking, Laser swap - updates 
                                  Reply  Fri Feb 12 17:05:40 2016, gautam, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                     Reply  Fri Feb 12 19:07:52 2016, Koji, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                        Reply  Fri Feb 19 00:42:19 2016, gautam, Update, Green Locking, Laser swap - beam ellipticity from laser? IMG_6287.JPGIMG_6288.JPGIMG_6286.JPG
                                           Reply  Mon Feb 22 13:56:52 2016, gautam, Update, Green Locking, Laser swap -reflected beam from ETM aligned 
                                              Reply  Tue Feb 23 23:01:16 2016, gautam, Update, Green Locking, Laser swap - Green PDH locking AM_scan.pdfAM_scan_zoomed.pdf
                                                 Reply  Wed Feb 24 19:29:13 2016, gautam, Update, Green Locking, Laser swap - Green PDH locked 
                                                    Reply  Fri Feb 26 01:52:44 2016, gautam, Update, Green Locking, Laser swap - Green PDH OLTF X_PDH_OLTF_20160225.pdf
                                                       Reply  Mon Feb 29 17:17:26 2016, gautam, Update, Green Locking, Laser swap - still no green beatnote 
                                                          Reply  Wed Mar 2 17:42:19 2016, gautam, Update, Green Locking, Laser swap - some progress 
                                                             Reply  Thu Mar 3 01:25:50 2016, gautam, Update, Green Locking, Laser swap - 2 IR + 1 green beatnotes found IR_beat_20160303_2.pdf
                                                                Reply  Fri Mar 4 01:11:41 2016, gautam, Update, Green Locking, Laser swap - both green beatnotes found IR_beat_20160303_green.pdf
                                                                   Reply  Sat Mar 5 23:31:01 2016, gautam, Update, Green Locking, Laser swap - some updates AUXxTelescope.png.pngPM_setup.pdfPLLolg.pdfAMPM20160303.pdfIRbeat_20160304.pdf
                                                                      Reply  Mon Mar 7 23:51:36 2016, gautam, Update, Green Locking, Laser swap - some improvement PomonaTF.pdfXPDH.pdfYPDH.pdfgreenbeat_20160307.pdf
                                                                         Reply  Tue Mar 8 18:22:20 2016, rana, Update, Green Locking, Laser swap - some improvement 
Message ID: 11984     Entry time: Tue Feb 9 19:15:36 2016     In reply to: 11983     Reply to this: 11985
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Laser swap - updates 

Some updates on the laser swap situation:

  1. Mode-matching calculation: 
  • I should have caught this earlier, but it was an oversight - the 35um waist that Andres used in his calculation is the waist size of the green beam. So I've been off by a factor of sqrt(2) all this while, and it works out that the desired waist size is indeed 50um, consistent with Kiwamu's elogs. Furthermore, as he has detailed in that elog, we actually want the free-space waist of the input beam to the doubling crystal to be ~6.7mm from the geometric center of the PPKPT crystal. 
  • I redid the calculation using these updated numbers. Attachment #1 shows the results (optimized for the X-waist, Y-profile plotted for comparison and to see what mode-matching efficiency we get). The way I've set up the code is for a la mode to rank the solutions in order of increasing sensitivity to the positions of the lenses. It turns out the least sensitive solution doesn't actually achieve the desired waist size of 50um - moreover, it requires us to change both lenses currently in the path. The next lease sensitive solution, however, achieves the desired waist (i.e. 100% theoretical mode-matching efficiency for the X mode) and only requires us to swap the 125mm lens we put in yesterday for a 150mm lens (and the positions of the lenses change slightly compared to what we had yesterday as well). The sensitivity in a la mode is parametrized by the amount of power remaining in the TEM00 mode while displacing one or more components. It turns out that this figure of merit is only ~1% smaller for the 2nd least sensitive solution compared to the first. So I've chosen to use that solution. Code used to calculate the mode matching is Attachment #2.
  • I've also plotted in Attachment #1 what the beam profile would have looked like before our modificatons last night, using the numbers from Zach's elog - as I have already mentioned in the previous elog, it suggests that the waist size would have been 39um, at a location 1.0821m in my coordinate system (desired position according to considerations in the previous 2 bullets is 1.0713). This seems to have been a sub-optimal configuration, but is also subject to errors I made in measuring the positions of the mirrors/lenses (I don't think I had 1cm resolution).

       2. Implementing the new solution:

  • I've switched out the 125mm efl lens for a 150mm efl lens from the same Thorlabs lens kit. I've also moved both the lenses to their new appropriate positions.
  • Unfortunately, I had put in some irides in the beampath before calculating this new (more appropriate solution). As a result, both the lenses are off from their optimal positions by a few mm because the irides get in the way. I guess we just have to live with this for now, and can adjust the positions of the lenses once we actually get some green light and are happy with all the other alignments...
  • As noted in the previous elog, I suspect that we saw no green light yesterday because we were missing the doubling crystal altogether (given that we have only a 1mm x 1mm area to aim for - the Faraday serves as a coarse constraint, though its aperture has ~25times this area!). I tried playing around with the two steering mirrors immediately after the NPRO to see if I could get some green light out, but have not been successful yet. I may make some further trials later in the evening/tomorrow...

As I check the manual of the Innolight (pg17) and the datasheet of the Lightwave, I wonder if the Quarter Wave Plate that was placed immediately after the Innolight laser head is even necessary now - I assume the purpose of the combination of QWP+HWP was to turn the elliptically polarized light from the Innolight into linearly polarized light before the Faraday. But the Lightwave already produces linearly polarized light. I will check out what is the configuration on the Y-end table...

 

Attachment 1: Modematch_X.pdf  176 kB  | Show | Show all
Attachment 2: XendModeMatch.m.zip  1 kB
ELOG V3.1.3-